Find the possible value or values of n in the quadratic equation 2n2 – 7n + 6 = 0.
Answer by stanbon(75887)

Find the possible value or values of n in the
quadratic equation 2n^2 – 7n + 6 = 0.
----------------------
Factor:
2n^2 - 3n - 4n + 6 = 0
n(2n-3) - 2(2n-3) = 0
(2n-3)(n-2) = 0
n = 3/2 or n = 2
----------------------
Cheers,
Stan H.